OpenStudy (anonymous):

I give medals!!!!!!! Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5 A. y – 4 = 5(x + 2) B. y – 4 = 5(x – 2) C. y + 4 = 5x – 10 D. y + 4 = 5(x + 2)

OpenStudy (anonymous):

do you know point slope form of line???

OpenStudy (anonymous):

no thats all they give me

OpenStudy (anonymous):

you are right but its all i m asking you

OpenStudy (anonymous):

do you know this y-y1 = m(x-x1)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its the point slope form of a line now you have the general equation y-y1 = m(x-x1) Point = (x1, y1) = (–2, 4) ; Slope= m = 5 put these values in the above equation and get your answer

OpenStudy (anonymous):

can you show me i dont get it

OpenStudy (anonymous):

ok you have this general form y-y1 = m(x-x1) and the given data is Point = (x1, y1) = (–2, 4) ; Slope= m = 5 y-4 =5(x-(-2)) do the last step now and get your answer

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

\[\large y-\color{red}{y_1}=\color{blue}m(x-\color{green}{x_1})\] put \(\color{red}{y_1}=\color{red}4\) \(\color{blue}{m}=\color{blue}5\) \(\color{green}{x_1}=\color{green}{-2}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-4=5(x-(-2))

OpenStudy (anonymous):

now -(-) = +

OpenStudy (anonymous):

y-54=5(x+2)

OpenStudy (anonymous):

y-4

OpenStudy (anonymous):

yap you are right so which is your best choice, A , B , C ,or D

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

exactly, you got it right

OpenStudy (anonymous):

i have one more plz

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3

OpenStudy (anonymous):

repeat the same process as i told you before

OpenStudy (anonymous):

y-(-7)(-2/3)(x-1)

OpenStudy (anonymous):

y-(-7) = (-2/3)(x-1) now apply the rule -(-) = +

OpenStudy (anonymous):

y+7=(-2/3)(x-1)

OpenStudy (anonymous):

so its your answer

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

welcome